OpenStudy (anonymous):

Simplify the expression 18 Squareroot 50 .

OpenStudy (anonymous):

\[18\sqrt{50}=90\sqrt{2}\]

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

\[y*\sqrt{x}=\sqrt{y^2x}\] find the y that satisfies the condition with a simplified x

OpenStudy (anonymous):

oh thank you so much this helps because i have EOCT tommorow